# Learning Versatile Humanoid Manipulation with Touch Dreaming

### 14 Apr 2026 v1

@@ Abstract
Humanoid robots promise general-purpose assistance, yet real-world humanoid loco-manipulation remains challenging because it requires whole-body stability, dexterous hands, and contact-aware perception under frequent contact changes. In this work, we study dexterous, contact-rich humanoid loco-manipulation. We first develop an RL-based whole-body controller that provides stable lower-body and torso execution during complex manipulation. Built on this controller, we develop a whole-body humanoid data collection system that combines VR-based teleoperation with human-to-humanoid motion mapping, enabling efficient collection of real-world demonstrations. We then propose Humanoid Transformer with Touch Dreaming (HTD), a multimodal encoder–decoder Transformer that models touch as a core modality alongside multi-view vision and proprioception. HTD is trained in a single stage with behavioral cloning augmented by touch dreaming: in addition to predicting action chunks, the policy predicts future hand-joint forces and future tactile latents, encouraging the shared Transformer trunk to learn contact-aware representations for dexterous interaction. Across five contact-rich tasks, Insert-T, Book Organization, Towel Folding, Cat Litter Scooping, and Tea Serving, HTD achieves a 90.9% relative improvement in average success rate over the stronger baseline. Ablation results further show that latent-space tactile prediction is more effective than raw tactile prediction, yielding a 30% relative gain in success rate. These results demonstrate that combining robust whole-body execution, scalable humanoid data collection, and predictive touch-centered learning enables versatile, high-dexterity humanoid manipulation in the real world. Project webpage: humanoid-touch-dream.github.io.
Figure 1: Our system enables versatile, contact-rich, and dexterous humanoid manipulation. A: long-horizon, multi-stage manipulation of deformable objects (towel folding). B: mixed prehensile and non-prehensile manipulation for thin-profile rigid objects with limited grasp affordance (book organization). C: tight-tolerance insertion with a clearance of 3.5 mm, requiring high precision and reactive adaptation (Insert-T). D: dexterous, tool-mediated contact under low-profile constraints (cat litter scooping). E: bimanual object fetch and loco-manipulation, requiring stable whole-body transport while keeping objects balanced and undisturbed (tea serving).
@@ I Introduction
Humanoid robots promise general-purpose physical assistance, fueled by rapid progress in whole-body control, teleoperation, and humanoid learning systems [14, 29, 44, 46, 11, 32]. Yet real-world humanoid loco-manipulation remains fundamentally challenging because it requires the tight coordination of whole-body stability, full end-effector dexterity, and contact-aware perception. In contact-rich tasks, small pose or force errors can quickly cascade into slip, jamming, or loss of balance. These challenges are especially acute for humanoids, where dexterous hand interaction is tightly coupled with torso posture, locomotion, and foot–ground support [25, 9]. As a result, accurate hand motion alone is not enough; successful humanoid manipulation also requires robust whole-body execution and timely understanding of contact.
@@
A first bottleneck is system capability. Contact-rich humanoid manipulation requires a practical real-world pipeline that jointly supports stable whole-body execution, full dexterous-hand control, and tactile sensing. Although recent humanoid systems have improved motion tracking, teleoperation, and demonstration collection [52, 28, 59], Table I highlights that few systems combine whole-body control, full end-effector dexterity, and touch sensing/modeling in a single platform for dexterous, contact-rich manipulation. To address this, we build an integrated whole-body humanoid manipulation system that combines an RL-based whole-body controller (WBC) with VR teleoperation, upper-body IK, dexterous hand retargeting, and distributed tactile sensing. This design provides a stable platform for collecting high-quality real-world demonstrations while allowing the operator to focus on task intent and dexterous interaction.
@@
A second bottleneck is representation learning. Purely action-supervised behavioral cloning from vision and proprioception often struggles in contact-rich manipulation because contact is only partially observed and can change abruptly [24]. Tactile sensing is therefore a natural complementary modality, and prior work has shown its value in visuo-tactile manipulation and predictive tactile learning [5, 17, 47, 58]. Yet most existing tactile learning methods are developed for arm-hand manipulation and often rely on separate tactile pretraining, explicit world-model modules, multi-stage inference, or manually designed virtual targets tied to specific tactile layouts [56, 47, 58, 49, 8]. More broadly, predictive latent learning in Joint-Embedding Predictive Architectures such as I-JEPA [1] and V-JEPA2 [2] suggests that future prediction in latent space can induce semantically meaningful representations without reconstructing raw observations or training a separate generative pipeline. However, these ideas have rarely been brought into a single-stage whole-body humanoid imitation policy that must jointly handle dexterous manipulation, locomotion-related action generation, and rapidly changing contact.
@@
We evaluate the full system on five real-world contact-rich tasks: Insert-T, Book Organization, Towel Folding, Cat Litter Scooping, and Tea Serving. These tasks span tight-tolerance insertion, hard-to-grasp rigid-object manipulation, long-horizon deformable-object handling, low-profile tool use, and bimanual loco-manipulation, together stressing precise alignment, sustained contact, whole-body coordination, and diverse interaction modes. Across these tasks, HTD achieves a 90.9% relative improvement in average success rate over the stronger ACT baseline, while ablations show that latent tactile dreaming is more effective than raw tactile prediction. Together, these results suggest that combining robust whole-body execution, integrated dexterous manipulation hardware, and predictive touch-centered learning is a practical path toward more reliable humanoid manipulation under frequent and complex contact changes.
@@
  • We develop a whole-body humanoid manipulation system that combines VR teleoperation with an RL-based whole-body controller for stable and accurate real-world humanoid manipulation.
  • We introduce Humanoid Transformer with Touch Dreaming (HTD), a multimodal encoder–decoder Transformer for humanoid loco-manipulation that models touch as a core modality and is trained in a single stage with touch dreaming, including future hand-joint-force prediction and EMA-supervised tactile-latent prediction for contact-aware representation learning.
  • We evaluate our full system on five real-world contact-rich humanoid manipulation tasks spanning insertion, rigid-object reorientation, deformable-object handling, tool use, and bimanual loco-manipulation, and show that HTD achieves strong gains over baselines, including a 90.9% relative improvement in average success rate.
@@ III-A A System for Versatile Humanoid Dexterous Manipulation
Fig. 2 presents our system for learning real-world, dexterous, contact-rich humanoid manipulation. The system consists of four stages: lower-body controller (LBC) training, VR-based teleoperation and data collection, policy learning with Humanoid Transformer with Touch Dreaming (HTD), and deployment. At its foundation is an RL-based LBC that provides stable lower-body and torso execution during manipulation. We train this controller in simulation with a teacher–student framework: a teacher policy learns robust lower-body behaviors under retargeted arm motions, and a deployable student policy imitates it using only proprioception and short history. The resulting LBC tracks base velocity, torso orientation, and height commands, and serves as the execution backbone during both teleoperation and deployment.
@@ III-B Lower-body Controller
Training details. During training, command signals are uniformly sampled within predefined ranges to cover diverse locomotion behaviors. We also apply domain randomization to improve sim-to-real transferability.
@@ III-C Teleoperation and Data Collection
These targets are executed through a three-stage stack. First, the LBC takes $(\bm{v},\bm{rpy},\bm{h})$ and produces lower-body joint targets $\bm{q}_{\mathrm{lower}}$ to maintain stable locomotion, posture, and torso tracking. Second, an IK solver maps the desired wrist/end-effector poses $\bm{x}_{\mathrm{wrist}}$ to upper-body joint targets $\bm{q}_{\mathrm{upper}}$. Third, a hand retargeting module based on DexPilot [12] converts the human hand targets $\bm{x}_{\mathrm{hand}}$ into dexterous hand joint targets by optimizing fingertip-distance consistency for reliable grasping and in-hand interaction. Together, this stack enables coordinated whole-body teleoperation while preserving full end-effector dexterity.
During teleoperation, we record synchronized multimodal observations from the humanoid, including RGB images from the dual-lens head camera and wrist cameras, robot and hand proprioception, per-joint force feedback from the dexterous hands, and tactile readings from both hands. Each hand provides a 1062-dimensional tactile observation distributed over 17 spatial sensing regions spanning the finger segments and palm surfaces. This distributed tactile layout captures localized contact patterns across the hand surface and is visualized in Fig. 3. The resulting dataset therefore pairs whole-body action targets with multi-view vision, robot and hand proprioception, per-joint hand-force feedback, and distributed tactile observations for downstream policy learning.
@@ IV Experiments
We conduct experiments to answer the following research questions: 1. How does our WBC strategy compare to other methods with regards to tracking accuracy and robustness? 2. How is the overall performance of HTD on real-world versatile humanoid manipulation? 3. How do touch and touch dreaming contribute to dexterous, contact-rich manipulation?
A) How does our WBC strategy compare to other methods with regards to tracking accuracy and robustness? To evaluate tracking performance, we benchmark our WBC against two leading humanoid locomotion controllers: FALCON [55]: employs a dual-policy architecture with separate control for upper and lower body, featuring adaptive force curriculum learning. AMO [25]: utilizes a hierarchical framework integrating trajectory optimization with RL-based motion adaptation, where legs and waist are regulated through a tracking controller while arms use PD control. We quantify performance with tracking error metrics:
@@
Our WBC achieves the best overall tracking on most metrics in Table II, notably linear velocity and torso height and orientation ($E_{h}$, $E_{y}$, $E_{p}$, $E_{r}$), indicating tighter posture regulation that is crucial for contact rich whole body manipulation. While AMO attains a slightly lower yaw rate error $E_{\omega}$, our controller provides a stronger balance between locomotion tracking and whole body configuration control. We note that the standard deviations of $E_{h}$ and especially $E_{p}$ exceed their means, largely due to a small number of difficult command combinations that create conflicting objectives, e.g., simultaneously requesting large forward torso pitch and a low base height; in these cases the policy may temporarily trade off tracking for stability, increasing variance despite low average errors across the command space.
@@
Average tracking error alone does not fully characterize the usable operating envelope of a whole body controller, since manipulation oriented control also requires stability under large torso reorientation and height variation that expand the reachable workspace. We therefore measure the per dimension stable controllable range of our WBC policy in simulation by sweeping one command dimension at a time, base height $h$, torso roll $\phi_{\text{torso}}$, torso pitch $\theta_{\text{torso}}$, and torso yaw $\psi_{\text{torso}}$, while keeping the others nominal, and running the policy until instability, a fall, or persistent tracking failure occurs; we report the largest interval that remains stable over the rollout. Our policy achieves stable ranges of $h\in[0.33,\,0.80]$ m, $\theta_{\text{torso}}\in[-0.92,\,1.41]$ rad, $\phi_{\text{torso}}\in[-0.38,\,0.35]$ rad, and $\psi_{\text{torso}}\in[-1.50,\,1.34]$ rad, covering the trained height range and most of the yaw range, with a broad but asymmetric pitch range and a comparatively narrower roll range that remains the most restrictive direction. Fig. 5 visualizes postures near the boundary, showing that our controller supports a wide stable region for crouching, bending, and large torso reorientation beyond nominal upright motions.
Figure 5: Visualization of postures near the boundary of the stable controllable workspace of our WBC policy in simulation.
B) How does our data collection and learning system for humanoid manipulation work in real-world settings? To answer this question, we evaluate our learning system on five contact-rich humanoid manipulation tasks:
@@
Baselines. We compare our approach against two decoder-only variants of ACT [57], which we found empirically perform better than the version augmented with a CVAE encoder. ACT (Visual + Proprio) uses only multi-view vision and proprioception. ACT (Visual + Proprio + Touch) additionally takes force and tactile observations as input. We report results for our method, HTD, which augments imitation learning with touch dreaming.
@@
C) How do touch and touch dreaming contribute to dexterous, contact-rich manipulation?
@@
(a) Tea Serving. Heatmaps correspond to the right middle finger.
(b) Towel Folding. Heatmaps correspond to the left index finger.
@@
We further isolate the contribution of touch observations and touch dreaming via ablations of HTD. Specifically, we evaluate four variants: w/o Touch and TD removes both touch observations and the touch dreaming objective. w/o TD keeps touch observations as input but removes the dreaming loss. Dream Raw Tactile predicts future tactile signals in the raw sensor space. Dream Latent Tactile (our full method) predicts future tactile signals in a learned latent space supervised by an EMA teacher. All variants are trained with the same behavioral cloning objective for action chunk prediction, differing only in touch-related inputs and auxiliary objectives.
@@
Qualitative touch dreaming. Fig. 8 illustrates the representations learned via the dreaming objective. Across representative rollouts, HTD predicts future hand force trajectories that effectively track both the timing and magnitude of contact events. Furthermore, tactile latent similarity (L2 similarity) remains high during periods of sustained contact. While we observe temporary drops in similarity during abrupt contact transitions, which coincide with sudden force spikes, the metric maintains a relatively high level overall. These localized deviations are expected: because we roll out the dreamed latent chunks in an open-loop manner, predictions naturally diverge slightly from the ground truth when unpredictable, discontinuous contact changes occur mid-chunk.
Comparing the two tasks further highlights the contact-awareness of the learned representations. In general, Tea Serving involves rigid objects and requires larger applied forces than the deformable manipulation in Towel Folding. During phases with light or sparse contact (e.g., Tea Serving at 6.7s, or the preliminary contacts in Towel Folding at 19.7s and 32.7s), the baseline latent patterns remain highly consistent across different fingers and tasks. Conversely, when rich contact occurs, the latents activate into distinct, high-intensity patterns (e.g., Tea Serving at 13.3s, 20.0s, and 26.7s; Towel Folding at 6.7s and 45.7s). Notably, when fingers experience comparable contact states, such as similar force magnitudes and localized contact regions, the resulting tactile latents exhibit visually analogous structural patterns. This consistency across varying levels of contact intensity suggests that our learned tactile latent space effectively filters out the high-frequency noise and spatial sparsity typical of raw sensor signals. Ultimately, these qualitative results reinforce our quantitative ablations: they demonstrate that the model has learned a robust, noise-filtered representation that accurately captures physical interaction, ensuring that the transformer trunk remains highly contact-aware during downstream manipulation.

### 27 Apr 2026 v2

@@ Abstract
Humanoid robots promise general-purpose assistance, yet real-world humanoid loco-manipulation remains challenging because it requires whole-body stability, end-effector dexterity, and contact-aware interaction under frequent contact changes. In this work, we study dexterous, contact-rich humanoid loco-manipulation. We first develop an RL-based lower-body controller that serves as the stability backbone for whole-body execution during complex manipulation. Built on this controller, we develop a VR-based whole-body humanoid data collection system that integrates dexterous hands and tactile sensing for contact-rich manipulation. We then propose Humanoid Transformer with Touch Dreaming (HTD), a multimodal encoder–decoder Transformer that models touch as a core modality alongside multi-view vision and proprioception. HTD is trained in a single stage with behavioral cloning augmented by touch dreaming: in addition to predicting action chunks, the policy predicts future hand-joint forces and future tactile latents, with tactile-latent targets provided by an exponential moving average target encoder without requiring a separate tactile pretraining stage. This encourages the policy to learn contact-aware representations for dexterous manipulation. Across five real-world contact-rich tasks, HTD achieves a 90.9% relative improvement in average success rate over the stronger baseline. Ablation results further show that latent-space tactile prediction is more effective than raw tactile prediction, yielding a 30% relative gain in success rate. These results demonstrate that our touch-dreaming-enhanced learning system enables versatile, high-dexterity humanoid manipulation in the real world. More information and open-source materials are available at humanoid-touch-dream.github.io.
Figure 1: Our system enables versatile, contact-rich, and dexterous humanoid manipulation. A: long-horizon, multi-stage manipulation of deformable objects (towel folding). B: mixed prehensile and non-prehensile manipulation for thin-profile rigid objects with limited grasp affordance (book organization). C: tight-tolerance insertion with a clearance of 3.5 mm, requiring high precision and reactive adaptation (Insert-T). D: dexterous, tool-mediated contact under low-profile constraints (cat litter scooping). E: bimanual object fetch and loco-manipulation, requiring stable whole-body transport while keeping objects balanced (tea serving); the embedded panel visualizes touch dreaming by showing the dreamed tactile latent of the right middle finger as a normalized heatmap that changes with finger-object contact.
@@ I Introduction
Humanoid robots promise general-purpose physical assistance, fueled by rapid progress in whole-body control, teleoperation, and humanoid learning systems [14, 29, 44, 46, 11, 32]. Yet real-world humanoid loco-manipulation remains fundamentally challenging because it requires tight coordination between whole-body stability and end-effector dexterity under frequent contact changes. Unlike fixed-base platforms, humanoid manipulation is physically coupled with torso posture, base motion, and foot-ground stability, so uncertain hand-object contact can affect not only local manipulation accuracy but also whole-body execution. Consequently, accurate hand motion alone is insufficient; reliable humanoid manipulation requires stable whole-body coordination together with contact-aware interaction.
@@
Despite these requirements, few existing systems provide a practical real-world pipeline that jointly supports stable whole-body execution, full dexterous-hand control, and tactile sensing. Although recent humanoid systems have improved motion tracking, teleoperation, and demonstration collection [52, 28, 59], Table I highlights that few systems combine whole-body control, full end-effector dexterity, and touch sensing/modeling in a single platform for dexterous, contact-rich manipulation. To address this, we build an integrated whole-body humanoid manipulation system that combines VR teleoperation with an RL-based lower-body controller (LBC), upper-body IK, dexterous hand retargeting, distributed tactile sensing, and touch modeling. This design provides a stable execution backbone for collecting high-quality real-world demonstrations while allowing the operator to focus on task intent and dexterous interaction.
@@
Built on this tactile-enabled system, the next challenge is how to learn contact-aware policies from multimodal humanoid demonstrations. Purely action-supervised behavioral cloning from vision and proprioception can struggle in contact-rich manipulation, where contact is partially observed and changes abruptly over time [24]. Tactile sensing provides a natural complementary signal for capturing these contact dynamics, and prior work has demonstrated its value in visuo-tactile manipulation and predictive tactile learning [5, 17, 47, 58]. Yet most existing tactile learning methods are developed for arm-hand manipulation and often rely on separate tactile pretraining, explicit world-model modules, multi-stage inference, or manually designed virtual targets tied to specific sensing and control setups [56, 47, 58, 49, 8]. More broadly, predictive latent learning in Joint-Embedding Predictive Architectures such as I-JEPA [1] and V-JEPA2 [2] suggests that future prediction in latent space can induce semantically meaningful representations without reconstructing raw observations or training a separate generative pipeline. However, these ideas have rarely been brought into a single-stage whole-body humanoid imitation policy that must jointly handle dexterous manipulation, locomotion-related action generation, and rapidly changing contact.
@@
We evaluate the full system on five real-world contact-rich tasks: Insert-T, Book Organization, Towel Folding, Cat Litter Scooping, and Tea Serving. These tasks span tight-tolerance insertion, hard-to-grasp rigid-object manipulation, long-horizon deformable-object handling, low-profile tool use, and bimanual loco-manipulation, together stressing precise alignment, sustained contact, whole-body coordination, and diverse interaction modes. Across these tasks, HTD achieves a 90.9% relative improvement in average success rate over the stronger ACT baseline, while ablations show that latent tactile dreaming is more effective than raw tactile prediction. Together, these results suggest that combining stable whole-body execution, tactile-enabled dexterous manipulation, and predictive touch dreaming provides a practical path toward reliable humanoid manipulation under frequent contact changes.
@@
  • We develop a tactile-enabled whole-body humanoid manipulation system for stable, dexterous, contact-rich real-world manipulation.
  • We introduce Humanoid Transformer with Touch Dreaming (HTD), a multimodal encoder–decoder Transformer for humanoid loco-manipulation that treats touch as a core modality and is trained in a single stage with touch dreaming, including future hand-joint-force prediction and EMA-supervised tactile-latent prediction for contact-aware representation learning.
  • We evaluate the full framework on five real-world contact-rich humanoid manipulation tasks and show that HTD consistently improves over baselines, with latent tactile prediction outperforming raw tactile prediction.
@@ III-A A System for Versatile Humanoid Dexterous Manipulation
Fig. 2 presents our system for learning real-world, dexterous, contact-rich humanoid manipulation. The system consists of four stages: lower-body controller (LBC) training, VR-based teleoperation and data collection, policy learning with Humanoid Transformer with Touch Dreaming (HTD), and deployment. At its foundation is an RL-based LBC that provides stable lower-body and torso execution during manipulation. We train this controller in simulation with a teacher–student framework: a teacher policy learns robust lower-body behaviors under retargeted arm motions, and a deployable student policy imitates it using only proprioception and short history. The resulting LBC tracks base velocity, torso orientation, and height commands, and serves as the execution backbone during both teleoperation and HTD policy deployment.
@@ III-B Lower-body Controller
Training details. During training, command signals are uniformly sampled from predefined ranges to cover diverse loco-manipulation behaviors. We also apply domain randomization to improve sim-to-real transferability. Additional details of LBC training are provided in Sec. V-A in Appendix.
@@ III-C Teleoperation and Data Collection
These targets are executed through a three-stage stack. First, the LBC takes $(\bm{v},\bm{rpy},\bm{h})$ and produces lower-body joint targets $\bm{q}_{\mathrm{lower}}$ to maintain stable locomotion, posture, and torso tracking. Second, an IK solver maps the desired wrist/end-effector poses $\bm{x}_{\mathrm{wrist}}$ to upper-body joint targets $\bm{q}_{\mathrm{upper}}$. Third, a hand retargeting module based on DexPilot [12] converts the human hand targets $\bm{x}_{\mathrm{hand}}$ into dexterous hand joint targets $\bm{q}_{\mathrm{hand}}$ by optimizing fingertip-distance consistency for reliable grasping and in-hand interaction. Together, this stack enables coordinated whole-body teleoperation while preserving full end-effector dexterity.
During teleoperation, we record synchronized multimodal observations from the humanoid, including RGB images from the dual-lens head camera and wrist cameras, robot and hand proprioception, per-joint force feedback from the dexterous hands, and tactile readings from both hands. Each hand provides a 1062-dimensional tactile observation distributed across 17 spatial sensing regions spanning the finger segments and palm surfaces. This distributed tactile layout captures localized contact patterns over the hand surface, as visualized in Fig. 3. The resulting dataset pairs whole-body action commands ($\bm{v},\bm{rpy},\bm{h}$, $\bm{x}_{\mathrm{wrist}}$, and $\bm{x}_{\mathrm{hand}}$) with multi-view vision, robot and hand proprioception, per-joint hand-force feedback, and distributed tactile observations for downstream policy learning.
@@ IV Experiments
We conduct experiments to answer the following research questions: 1. How does our LBC strategy compare to other methods with regards to tracking accuracy and robustness? 2. How do our learning system and HTD perform on real-world versatile humanoid manipulation? 3. How do predictive touch dreaming and latent tactile supervision impact the effectiveness of contact-rich humanoid manipulation? 4. Does touch dreaming capture meaningful and robust contact-aware representations?
IV-A How does our LBC strategy compare to other methods with regards to tracking accuracy and robustness?
To evaluate tracking performance, we benchmark our LBC against two representative decoupled whole-body humanoid controllers: FALCON [55], which uses a dual-policy architecture with separate upper- and lower-body control and adaptive force curriculum learning; and AMO [25], which uses a hierarchical whole-body control framework where upper-body goals and torso commands are converted into lower-body references through an adaptive motion optimization module and tracked by a learned lower policy. We quantify performance with tracking error metrics:
@@ IV-A How does our LBC strategy compare to other methods with regards to tracking accuracy and robustness?
Figure 5: Visualization of postures near the boundary of the stable controllable workspace of our LBC policy in simulation.
Our LBC achieves the best overall tracking on most metrics in Table II, notably linear velocity and torso height and orientation ($E_{h}$, $E_{y}$, $E_{p}$, $E_{r}$), indicating tighter torso pose regulation that is crucial for contact-rich whole-body manipulation. While AMO attains a slightly lower yaw rate error $E_{\omega}$, our controller provides a stronger balance between locomotion tracking and whole-body configuration control. We note that the standard deviations of $E_{h}$ and especially $E_{p}$ exceed their means, largely due to a small number of difficult command combinations that create conflicting objectives, e.g., simultaneously requesting large forward torso pitch and a low base height; in these cases the policy may temporarily trade off tracking for stability, increasing variance despite low average errors across the command space.
@@
Average tracking error alone does not fully characterize the usable operating envelope of a whole-body controller, since manipulation-oriented control requires stability under large torso reorientation and height variation that expand reachability. We therefore measure the per-dimension stable controllable range of our LBC policy in simulation by sweeping one command dimension at a time, base height $h$, torso roll $\phi_{\text{torso}}$, torso pitch $\theta_{\text{torso}}$, and torso yaw $\psi_{\text{torso}}$, while keeping the others nominal, and running the policy until instability, a fall, or persistent tracking failure occurs; we report the largest interval that remains stable over the rollout. Our policy achieves stable command-tracking ranges of $h\in[0.33,\,0.80]$ m, $\phi_{\text{torso}}\in[-0.38,\,0.35]$ rad, $\theta_{\text{torso}}\in[-0.92,\,1.41]$ rad, and $\psi_{\text{torso}}\in[-1.50,\,1.34]$ rad. The stable workspace covers the trained height range and most of the yaw range, while roll remains most restrictive. Fig. 5 visualizes boundary postures, showing that our controller supports a wide stable region for crouching, bending, and large torso reorientation beyond nominal upright motions.
IV-B How do our learning system and HTD perform on real-world versatile humanoid manipulation?
To answer this question, we evaluate our learning system on five contact-rich humanoid manipulation tasks:
@@ IV-B How do our learning system and HTD perform on real-world versatile humanoid manipulation?
Baselines. We compare our approach against two decoder-only variants of ACT [57], which we found empirically perform better than the version augmented with a CVAE encoder. ACT (Visual + Proprio) uses only multi-view vision and proprioception. ACT (Visual + Proprio + Touch) additionally takes force and tactile observations as input. We report results for our method, HTD, which augments imitation learning with touch dreaming and employs a modular design for observations, action experts, and dream experts.
@@
IV-C How do predictive touch dreaming and latent tactile supervision impact the effectiveness of contact-rich humanoid manipulation?
@@
(a) Tea Serving. Heatmaps with raw latent values correspond to the right middle finger.
(b) Towel Folding. Heatmaps with raw latent values correspond to the left index finger.
@@
To answer this question, we further isolate the contribution of touch observations and touch dreaming via ablations of HTD. Specifically, we evaluate four variants: w/o Touch and TD removes both touch observations and the touch dreaming objective. w/o TD keeps touch observations as input but removes the dreaming loss. Dream Raw Tactile predicts future tactile signals in the raw sensor space. Dream Latent Tactile (our full method) predicts future tactile signals in a learned latent space supervised by an EMA teacher. All variants are trained with the same behavioral cloning objective for action chunk prediction, differing only in touch-related inputs and auxiliary objectives.
@@
IV-D Does touch dreaming capture meaningful and robust contact-aware representations?
Qualitative touch dreaming. To better understand what the touch-dreaming objective learns, we qualitatively examine whether the predicted future forces and tactile latents reflect meaningful contact dynamics during real-world rollouts. Fig. 8 illustrates the dreamed touch representations across two policy rollout episodes, Tea Serving and Towel Folding. Across these representative rollouts, HTD predicts future hand force trajectories that effectively track both the timing and magnitude of contact events. Furthermore, tactile latent similarity (L2 similarity) remains high during periods of sustained contact. On the right side of Fig. 8, we show the synchronized right-eye camera view together with heatmaps of the raw tactile latent values for the selected fingers at the timesteps indicated by the dashed vertical lines in the plots on the left. We observe that the predicted heatmaps are close to the ground-truth heatmaps at most timesteps. While temporary drops in similarity occur during abrupt contact transitions and coincide with sudden force spikes, the similarity remains relatively high overall. The localized deviations are expected for two reasons. First, because the dreamed latent chunks are rolled out in an open-loop manner, the predictions naturally diverge slightly from the ground truth when unpredictable, discontinuous contact changes occur mid-chunk. Second, the ground-truth tactile latents typically vary with the noisy raw tactile readings, whereas the dreamed tactile latents are more stable and more aligned with semantically meaningful contact changes inferred jointly from vision and applied forces, rather than directly matching raw tactile readings that can be noisy or weakly responsive.
Comparing the two tasks further highlights the contact-awareness of the learned representations. In general, Tea Serving involves rigid objects and requires larger applied forces than the deformable interaction in Towel Folding. During phases with light or sparse contact, such as Tea Serving at 6.7s or the preliminary contacts in Towel Folding at 19.7s and 32.7s, the latent patterns remain highly consistent across different fingers and tasks. Conversely, when richer contact occurs, the latents activate into distinct, high-intensity patterns, as seen in Tea Serving at 13.3s, 20.0s, and 26.7s, and in Towel Folding at 6.7s and 45.7s. Notably, when fingers experience comparable contact states, such as similar force magnitudes and localized contact regions, the resulting tactile latents exhibit visually analogous structural patterns. Meanwhile, the predicted tactile latents also show larger peak values under stronger contact and larger applied forces, for example when comparing the active contact patterns in Tea Serving with those in Towel Folding. This consistency across varying contact intensities suggests that the learned tactile latent space captures contact-relevant structure while being less affected by the spatial sparsity and local irregularity typical of raw sensor signals. Overall, these qualitative results reinforce the quantitative ablations by showing that the model learns a robust and contact-aware latent representation of physical interaction.